# Image Steganography For Securing Intellicise Wireless Networks: “Invisible Encryption” Against Eavesdroppers

### v1

@@ Abstract
Index Terms:
Intellicise wireless networks, semantic communications, image steganography, anti-eavesdropping.
@@ I Introduction
In addition to the aforementioned defense methods, image steganography [7, 8, 9] offers a promising approach for securing SemCom by embedding secret images into cover images, effectively concealing encryption traces. This technique, often referred to as “invisible encryption,” holds significant potential for enhancing the security of SemCom systems. By integrating steganography, the proposed secure SemCom approach by Tang et al. [10] effectively safeguards private image data during transmission, leveraging Invertible Neural Networks (INNs) to embed private images within host images. The method ensures high reconstruction quality for authorized receivers while denying eavesdroppers access to sensitive content. Building on this, Tang et al. [11] introduced novel intelligent eavesdroppers and demonstrated that, the INN-based steganography module can reduce eavesdropping success rates to 0%, while maintaining data quality for legitimate semantic receivers. Additionally, Huo et al. [12] employed Generative Adversarial Networks (GANs) to embed secret semantic features into cover semantic features, ensuring reliable extraction and maintaining covertness throughout SemCom. This approach not only preserves the integrity of SemCom but also strengthens security by addressing distortions at both pixel and semantic levels. Despite these advancements, research on image steganography-based SemCom is still in its nascent stages. The full potential of leveraging image steganography techniques to secure SemCom has yet to be thoroughly explored. Several key questions remain unanswered:
@@
Cryptography- based SemCom Traditional Cryptography Deep joint source-channel and encryption coding: Secure semantic communications: Based on learning with errors, and using the affine property of encryption method to realize privacy security. Lightweight and robust wireless semantic communications: AES encryption algorithm, adversarial identification, and SemCom are combined to resist eavesdropping attacks and adversarial attacks.
@@
Quantum Cryptography Quantum semantic communications for metaverse: Principles and challenges: The quantum semantic layer can ensure the privacy of semantic information through public keys generated by quantum key distribution. Cooperative resource management in quantum key distribution (QKD) networks for semantic communication: Analysis of the cooperation of quantum key distribution service providers. Integrating sustainable big AI: Quantum anonymous semantic broad-cast: Semantic extraction is realized by large models, and semantic information is broadcast by quantum anonymous communication protocol.
Covert SemCom Multi-agent reinforcement learning for covert semantic communications over wireless networks: Jointly maximizing semantic information transmission and power control to ensure the quality and security. Unveiling covert semantics: Joint source-channel coding under a covertness constraint: Veryfing the feasibility of training neural network to realize covert SemCom for classification tasks.
@@
Application- Layer Encrypted SemCom Cross Layer Encryption IRS-enhanced secure semantic communication networks: Cross-layer and context-awared resource allocation: Definitions of SSR and SSSE to quantify the semantic security performance of task level. Eavesdropping on semantic communication: Timing attacks and countermeasures: This study found that the privacy protection of SemCom is time-sensitive, and provides theoretical guidance based on Markov.
Neural Network- based Encryption Deep joint source-channel coding for image transmission with visual protection: Transforming the original image into a visually protected image. The model inversion eavesdropping attack in semantic communication systems: Introductions of the model inversion eavesdropping attack to reveal the risk of privacy leakage in SemCom system. Wireless image transmission with semantic and security awareness: An autoencoder based on residual module and convolution level is proposed to flexibly adjust the trade-off between efficiency and privacy. Disentangled information bottleneck guided privacy-protective joint source and channel coding for image transmission: A privacy-preserving SemCom method based on decoupling information bottleneck.
@@ III-A1 Convolutional Neural Network (CNN)-based Image Steganography JSCC
In traditional fully connected neural networks, each neuron connects to all inputs, leading to an explosion in the number of parameters. In contrast, neurons in CNNs only connect to local regions of the input data, imitating the local perception characteristics of human eyes, which greatly reduces the parameter count. Additionally, the same convolutional kernel slides across the entire input data to extract consistent types of features, such as edges and textures. This parameter-sharing mechanism further minimizes parameters while enhancing the model’s generalization capability. CNN-based image steganography JSCC typically include the following modules:
@@
  • At the transmitter, the CNN-based Semantic Steganography module covertly embeds the secret semantic features into the cover semantic features, leveraging CNN’s dynamic perception capabilities to select optimal embedding locations, resulting in stego semantic features. Thus, stego semantic features are the cover semantic features where secret semantic information is embedded. At the receiver, the CNN-based Semantic Steganography module extracts the reconstructed cover semantic features and reconstructed secret semantic features from the noisy stego semantic features.
@@
  • The Semantic Decoder restores the reconstructed cover semantic features and reconstructed secret semantic features into the reconstructed cover image and reconstructed secret image, respectively.
@@ III-A2 Generative Adversarial Network (GAN)-based Image Steganography JSCC
GAN is a deep learning model that employs adversarial training to pit two neural networks, the generator and the discriminator, against each other. Its objective is to generate synthetic data that closely resembles the distribution of real data, a minimax game $\min_{G}\max_{D}V(D,G)$. In this framework: the generator learns to create realistic data from random noise, such as Gaussian distributions, aiming to “deceive” the discriminator ($D(G(z))\to 1$), through minimizing $V(D,G)$; the discriminator acts as a classifier, striving to accurately determine whether input data is a real sample ($D(x)\to 1$) or a fake sample synthesized by the generator ($D(G(z))\to 0$), through maximizing $V(D,G)$. This adversarial interplay drives continuous improvement in both components, enabling the generator to produce highly convincing synthetic outputs. Therefore, in GAN-based image steganography JSCC, the generator is designed to produce stego semantic features, while the discriminator distinguishes between the cover semantic features encoded by the semantic encoder and the noisy stego semantic features. This adversarial process forces the generator to produce undetectable embeddings. Additionally, by modeling the presence of an intelligent eavesdropper via the discriminator’s detection attempts, this method enhances resistance to semantic steganalysis.
For instance, Huo et al. [12] used two discriminators to enhance both security and semantic extraction accuracy. The first discriminator focuses on distinguishing between cover and stego semantic features, prompting the generator to produce stego features that closely resemble the original cover features. The second discriminator operates after the semantic reconstruction stage, ensuring that even after image reconstruction, the embedded secret information remains secure and the system is resistant to steganalysis. In short, the first discriminator ensures that the stego features are undetectable during semantic transmission, while the second discriminator guarantees that secret information remains hidden even in the reconstructed image, thus enhancing the overall robustness and security of SemCom.
@@ III-A3 Invertible Neural Network (INN)-based Image Steganography JSCC
For instance, Tang et al. [10] introduced an INN-based signal steganography framework to enhance security in SemCom. The method leverages multiple invertible blocks and additive affine transformations to embed secret signals while preserving reversibility. The loss function incorporates both forward and backward propagation processes, along with a privacy protection term to safeguard sensitive information. Notably, the training process assumes no prior knowledge of Eves, i.e., an eavesdropper. Experimental results demonstrate that eavesdroppers employing naive decoding or model inversion attacks can only reconstruct cover images, validating the framework’s effectiveness against intelligent eavesdroppers. Tang et al. [11] conducted further assessments of the framework’s security performance against various intelligent eavesdropping scenarios, including glass-box, closed-box, GAI+Glass-box, and GAI+closed-box eavesdropping. Simulation results reveal that the proposed scheme effectively mitigates the risks associated with these eavesdropping threats. Additionally, to degrade the performance of eavesdroppers, Li et al. [13] proposed a hiding and deception method based on INNs. This method deceives eavesdroppers by generating virtual semantic information for a certain modality while simultaneously hiding the original semantic information of that modality into another modality. Legitimate users can extract the hidden semantic information and accomplish their tasks normally through INNs.
@@ III-B3 Self Learning-based Semantic Steganography
The self learning-based semantic steganography strategy leverages the end-to-end optimization capability of deep learning models to automatically learn the optimal way of embedding secret semantic features into cover semantic features without relying on manually predefined rules. Some typical models are as follows. AlexNet has a relatively small number of parameters and a simple structure, making it suitable for low-capacity semantic steganography in low-resolution images. VGGNet employs deep convolutional layers to capture more complex semantic features, enhancing the concealment of semantic feature embedding. GoogleNet utilizes Inception modules to simultaneously capture local details and global context, providing adaptive semantic region embedding. ResNet introduces residual structures, which extend the depth of networks and reduce the reconstruction error of secret semantic features. The attention mechanism can dynamically learn importance weights of different channels, thereby optimizing the deep semantic representation of images and improving the quality of stego semantic features as well as the accuracy of secret semantic feature extraction.
@@ III-B5 Pre-trained Model-based Semantic Steganography
Compared to single-stage end-to-end training, the pre-trained model-based semantic steganography strategy adopts a two-stage separable training approach. The semantic encoder and decoder focus on efficiently compressing and reconstructing semantic information, while the semantic steganography module concentrates on secret semantic feature embedding and extraction. This separable training strategy avoids direct conflicts between the objectives of the two types of tasks, preventing the image steganography JSCC model from falling into local optima. For example, Tang et al. [11] employed the DeepJSCC architecture [^1] https://arxiv.org/abs/1809.01733 as the trained semantic encoder and decoder, then froze their parameters, and further trained the parameters of INN-based signal steganography modules.
@@ V-A Theoretical Analysis of Semantic Steganography Capacity
According to the mathematical theory of SemCom [^4] https://arxiv.org/abs/2401.13387, we can know that, for band-limited Gaussian channels, the channel capacity can be represented as $C_{S}=\begin{aligned} B\log\left[S^{4}\left(1+\frac{P}{N_{0}B}\right)\right]\end{aligned}$, where $S$ is the average synonymous length. In steganographic SemCom, when we embed secret semantic features into cover semantic features, the effective channel capacity will decrease compared to SemCom without steganography. This is because, to achieve the steganographic goal and deceive semantic steganalyzers, we need to transmit cover information as well as the correlation information between cover and secret semantic features. To maximally confuse attackers, reducing the transmission rate of secret semantic features is a direct and effective method. However, doing so will inevitably degrade communication efficiency. Therefore, how to provide a theoretical analysis of semantic steganographic capacity based on channel capacity and steganographic capacity is crucial for guiding further research in the future.
@@ V-D Combination of Encryption and Steganography for Securing SemCom
Semantic steganography and semantic steganalysis are always in a state of confrontation, continuously evolving through this dynamic interplay. To provide more reliable privacy protection for SemCom, integrating encryption technology into semantic steganography is a promising solution. Homomorphic encryption allows direct homomorphic operations on encrypted data, ensuring that the privacy of the plaintext is not compromised during processing and providing security protection during semantic extraction and recovery [^5] https://arxiv.org/abs/2501.10182. Based on this, we can first perform semantic extraction on the ciphertext processed by homomorphic encryption. Then, we can design a semantic steganography module that supports homomorphic encryption to embed the secret semantic ciphertext into the cover semantic features. In this way, even if intelligent attackers obtain the secret semantic ciphertext through advanced semantic steganalysis techniques, they still need to crack the homomorphic encryption key to access the plaintext content.

### v2

@@ I Introduction
In addition to the aforementioned defense methods, image steganography [7, 3, 9] offers a promising approach for securing SemCom by embedding secret images into cover images, effectively concealing encryption traces. This technique, often referred to as “invisible encryption,” holds significant potential for enhancing the security of SemCom systems. Invertible Neural Networks (INNs) [11, 10, 5] and Generative Adversarial Networks (GANs) [4] have been employed to embed secret semantic features into cover semantic features, ensuring reliable extraction and maintaining covertness throughout SemCom. Nevertheless, research on image steganography-based SemCom is still in its nascent stages. The full potential of leveraging image steganography techniques to secure SemCom has yet to be thoroughly explored. Several key questions remain unanswered:
@@
Cryptography- based SemCom Traditional Cryptography Deep joint source-channel and encryption coding: Secure semantic communications: Based on learning with errors, and using the affine property of encryption method to realize privacy security.
@@
Quantum Cryptography Quantum semantic communications for metaverse: Principles and challenges: The quantum semantic layer can ensure the privacy of semantic information through public keys generated by quantum key distribution.
Covert SemCom Multi-agent reinforcement learning for covert semantic communications over wireless networks: Jointly maximizing semantic information transmission and power control to ensure the quality and security.
@@
Application- Layer Encrypted SemCom Cross Layer Encryption IRS-enhanced secure semantic communication networks: Cross-layer and context-awared resource allocation: Definitions of SSR and SSSE to quantify the semantic security performance of task level.
Neural Network- based Encryption The model inversion eavesdropping attack in semantic communication systems: Introductions of the model inversion eavesdropping attack to reveal the risk of privacy leakage in SemCom system.
@@ III-A1 Convolutional Neural Network (CNN)-based Image Steganography JSCC
Compared with the traditional fully connected neural network, the parameter sharing mechanism in CNN reduces the number of parameter, and uses convolution kernels to extract consistent features such as edges and textures, which can be used for image steganography SemCom. CNN-based image steganography JSCC typically include the following modules:
@@
  • At the transmitter, the CNN-based Semantic Steganography module covertly embeds the secret semantic features into the cover semantic features, leveraging CNN’s dynamic perception capabilities to select optimal embedding locations, resulting in stego semantic features. At the receiver, the CNN-based Semantic Steganography module extracts the reconstructed cover and secret semantic features from noisy stego semantic features.
@@
  • The Semantic Decoder restores the reconstructed cover and secret semantic features into the reconstructed cover and secret images, respectively.
@@ III-A2 Generative Adversarial Network (GAN)-based Image Steganography JSCC
GAN is a DL model that employs adversarial training to make two neural networks, the generator and the discriminator, against each other. Its objective is to generate synthetic data that closely resembles the distribution of real data, a minimax game $\min_{G}\max_{D}V(D,G)$. In this framework: the generator learns to create realistic data from random noise, such as Gaussian distributions, aiming to “deceive” the discriminator ($D(G(z))\to 1$), through minimizing $V(D,G)$; the discriminator acts as a classifier, striving to accurately determine whether input data is a real sample ($D(x)\to 1$) or a fake sample synthesized by the generator ($D(G(z))\to 0$), through maximizing $V(D,G)$. This adversarial interplay drives continuous improvement in both components, enabling the generator to produce highly convincing synthetic outputs. Therefore, in GAN-based image steganography JSCC, the generator is designed to produce stego semantic features, while the discriminator distinguishes between the cover semantic features encoded by the semantic encoder and the noisy stego semantic features. This adversarial process forces the generator to produce undetectable embeddings. Additionally, by modeling the presence of an intelligent eavesdropper via the discriminator’s detection attempts, this method enhances resistance to semantic steganalysis.
For instance, Huo et al. [4] used two discriminators to enhance both security and semantic extraction accuracy. The first discriminator ensures that the stego features are undetectable during semantic transmission, while the second discriminator guarantees that secret information remains hidden even in the reconstructed image, thus enhancing the overall robustness and security of SemCom.
@@ III-A3 Invertible Neural Network (INN)-based Image Steganography JSCC
For instance, Tang et al. [11] introduced an INN-based signal steganography framework to enhance security in SemCom. The method leverages multiple invertible blocks and additive affine transformations to embed secret signals while preserving reversibility. The loss function incorporates both forward and backward propagation processes, along with a privacy protection term to safeguard sensitive information. Experimental results demonstrate that eavesdroppers employing naive decoding or model inversion attacks can only reconstruct cover images, validating the framework’s effectiveness against intelligent eavesdroppers. Tang et al. [10] conducted further assessments of the framework’s security performance against various intelligent eavesdropping scenarios, including glass-box, closed-box, GAI+Glass-box, and GAI+closed-box eavesdropping. Simulation results reveal that the proposed scheme effectively mitigates the risks associated with these eavesdropping threats. Additionally, Li et al. [5] proposed a hiding and deception method based on INNs. This method deceives eavesdroppers by generating virtual semantic information for a certain modality while simultaneously hiding the original semantic information of that modality into another modality. Legitimate users can extract the hidden semantic information and accomplish their tasks normally through INNs.
@@ III-B3 Self Learning-based Semantic Steganography
The self learning-based semantic steganography strategy leverages the end-to-end optimization capability of deep learning models to automatically learn the optimal way of embedding secret semantic features into cover semantic features without relying on manually predefined rules. Some typical models include AlexNet, VGGNet, GoogleNet, ResNet, etc. In addition, the attention mechanism can dynamically learn importance weights of different channels, thereby optimizing the deep semantic representation of images and improving the quality of stego semantic features as well as the accuracy of secret semantic feature extraction.
@@ III-B5 Pre-trained Model-based Semantic Steganography
Compared to single-stage end-to-end training, the pre-trained model-based semantic steganography strategy adopts a two-stage separable training approach. The semantic encoder and decoder focus on efficiently compressing and reconstructing semantic information, while the semantic steganography module concentrates on secret semantic feature embedding and extraction. This separable training strategy avoids direct conflicts between the objectives of the two types of tasks, preventing the image steganography JSCC model from falling into local optima. For example, Tang et al. [10] employed the DeepJSCC architecture as the trained semantic encoder and decoder, then froze their parameters, and further trained the parameters of INN-based signal steganography modules.
@@ V-A Theoretical Analysis of Semantic Steganography Capacity
The mathematical theory of SemCom provides the expression of the channel capacity for band-limited Gaussian channels. In steganographic SemCom, when we embed secret semantic features into cover semantic features, the effective channel capacity will decrease compared to SemCom without steganography. This is because, to achieve the steganographic goal and deceive semantic steganalyzers, we need to transmit cover information as well as the correlation information between cover and secret semantic features. To maximally confuse attackers, reducing the transmission rate of secret semantic features is a direct and effective method. However, doing so will inevitably degrade communication efficiency. Therefore, how to provide a theoretical analysis of semantic steganographic capacity based on channel capacity and steganographic capacity is crucial for guiding further research in the future.
@@ V-D Combination of Encryption and Steganography for Securing SemCom
Semantic steganography and semantic steganalysis are always in a state of confrontation, continuously evolving through this dynamic interplay. To provide more reliable privacy protection for SemCom, integrating encryption technology into semantic steganography is a promising solution. Homomorphic encryption allows direct homomorphic operations on encrypted data, ensuring that the privacy of the plaintext is not compromised during processing and providing security protection during semantic extraction and recovery. Based on this, we can first perform semantic extraction on the ciphertext processed by homomorphic encryption. Then, we can design a semantic steganography module that supports homomorphic encryption to embed the secret semantic ciphertext into the cover semantic features. In this way, even if intelligent attackers obtain the secret semantic ciphertext through advanced semantic steganalysis techniques, they still need to crack the homomorphic encryption key to access the plaintext content.